OpenStudy (anonymous):

simplify 5* 4a + 6(5a +2) using distributive property

OpenStudy (swissgirl):

=5*4+6*(5a)+6*(2) =5*4+30a+12 =20+30a+12 =32+30a

OpenStudy (anonymous):

thanks!